# Super approximation for $\text{SL}_{2}\times\text{SL}_{2}$ and $\text{ASL}_{2}$

### v3

@@ 5.4. Creating a line
$M^{-1}[\sigma_{C_{3}},\cdots[\sigma_{2},[\sigma_{1},\zeta]]\cdots]M\equiv 1+\tilde{q}^{C_{3}}\left(\prod_{i=1}^{C_{3}}h_{\sigma_{i}}\right)X(\operatorname{mod}(\mathfrak{q}_{1}^{C_{3}}\mathfrak{q}_{2}^{2})$ (5.47)
where $X=\left(\begin{matrix}0&\zeta_{12}\\ (-1)^{C_{3}}\zeta_{21}&0\end{matrix}\right)$.
The sum structure is clear. For any $\sigma_{1,1},\cdots,\sigma_{C_{2},C_{3}}\in A_{6}$, we have
$(M^{-1}[\sigma_{1,C_{3}},\cdots[\sigma_{1,2},[\sigma_{1,1},\zeta]]M)\cdots(M^{-1}[\sigma_{C_{2},C_{3}},\cdots[\sigma_{C_{2,2}},[\sigma_{C_{{2},1}},\zeta]]\cdots]M)$
$\equiv$ $1+\tilde{q}^{C_{3}}\sum_{i=1}^{C_{2}}\left(\prod_{j=1}^{C_{3}}h_{\sigma_{i,j}}\right)X(\operatorname{mod}(\mathfrak{q}_{1}^{C_{3}}\mathfrak{q}_{2}^{2})$ (5.48)
From (5.33) and (5.4), we conclude there is a constant $K_{1}$ depending only on $c_{1}$, such that
$1+\tilde{q}^{C_{3}}\mathbb{Z}\xi X\subset M^{-1}A_{1}^{K_{1}}M(\operatorname{mod}\mathfrak{q}_{1}^{C_{3}}\mathfrak{q}_{2}\langle\mathfrak{a}^{\prime},\mathfrak{q}_{2}\rangle)$ (5.49)
Conjugating (5.49) back by $M$, noticing that $A_{1}^{K_{1}}$ is real and also having (5.35), (5.42), and (5.49) in mind, if we take
$\rho_{1}<\frac{c_{1}c_{3}}{28*192C_{1}}\rho_{2},$ (5.50)
then we obtain
@@
There exists $q_{3}||q_{1}$, $q_{3}>q^{\frac{c_{1}}{8C_{1}}}$, $q_{3}^{\prime}|q_{3}^{\prime\prime}|q_{3}$, some $K_{1}=K_{1}(c_{1})>0$, a matrix $X_{0}\in{{\text{Lie(SL}}}_{2})(\mathbb{Z})$ coprime to $q_{3}$, such that
$1+q_{3}^{\prime}\mathbb{Z}X_{0}\subset A_{1}^{K_{1}}(\operatorname{mod}q_{3}^{\prime\prime})$ (5.52)
where for each $p^{n}||q_{3}$, $v_{p}(q_{3}^{\prime\prime})=O_{c_{1}}(\rho_{2}n)$ and $\frac{c_{1}c_{3}}{192C_{1}}\rho_{2}n<v_{p}(q_{3}^{\prime\prime}/q_{3}^{\prime})<{v_{p}(q_{3}^{\prime\prime})}/2$.
5.5. Creating a segment
Next, we want to find $g_{1},g_{2}$ from a bounded product of $A_{1}$ so that $g_{1},g_{2}$ conjugates $X_{0}$ to different directions.
@@ 5.5. Creating a segment
If $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$, apply Proposition 4.2 with $\xi_{1}=X_{0},\xi_{2}=\eta_{2}=0$, and a proper choice of $\eta_{1}$ so that $\text{Tr}(g\xi_{1}g^{-1}\eta_{1})=L_{1}(g)$. If $\Lambda=\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$, apply Proposition 4.4 with $L=L_{1}$. Let $c_{2}$ be the implied constant from Propositions 4.2 and 4.4. Then we obtain, for any $Q>q_{3}^{\rho_{1}}$, we have
$\nu^{(l)}(\{g\in\text{SL}_{2}(\mathbb{Z})|L_{1}\circ T_{1}(g\tilde{X}g^{-1})\equiv 0(\operatorname{mod}{Q})\})<Q^{-c_{2}},$
where $l$ is as given in Proposition 2.2. This implies, arguing the same as for the construction of $A_{3}$, that there is $q_{4}||q_{3}$, $q_{4}>(q_{3})^{\frac{1}{2}}$, and $g_{1}\in A_{1}$ such that for any $p^{n}||q_{5}$, $L_{1}\circ T_{1}(g_{1}X_{0}g_{1}^{-1})\neq 0(p^{[2\rho_{1}n]})$, if we take
$\delta<\frac{c_{1}c_{2}\rho_{1}}{48C_{1}}.$ (5.53)
Let $X_{1}=g_{1}X_{0}g_{1}^{-1}$, then $X_{1}\equiv a(1,0,0)+b\vec{v}(\operatorname{mod}q_{5})$, where $a,b\in\mathbb{Z}$, and $b$ satisfies, for any $p^{n}||q_{5}$, $p^{[2\rho_{1}n]}\nmid b$, and $\vec{v}$ is a primitive vector in the span of $(0,1,0)$ and $(0,0,1)$.
Next, we choose an element $T_{2}\in\text{SL}_{3}(\mathbb{Z})$ such that $T_{2}(X_{0})=(1,0,0)$ and $T_{2}(X_{0})=(0,b,0)$. Let $L_{2}:\mathbb{Z}^{3}\mapsto\mathbb{Z}$ be the linear form getting the third component. Then $L_{2}\circ T_{2}$ is a primitive linear form on $\mathbb{Z}^{3}$. Applying Proposition 4.2 again, for any $Q>q_{4}^{\rho_{1}}$, we have
$\nu^{(l)}(\{g\in\text{SL}_{2}(\mathbb{Z})|L_{2}\circ T_{2}(g\tilde{X}g^{-1})\equiv 0(\operatorname{mod}{Q})\})<Q^{-c_{2}},$
which then implies there is $q_{5}||q_{4}$, $q_{5}>(q_{4})^{\frac{1}{2}}$, and $g_{2}\in A^{l}$ such that for any $p^{n}||q_{6}$, $L_{2}\circ T_{2}(g_{2}X_{0}g_{2}^{-1})\neq 0(p^{[2\rho_{1}n]})$, if
${\delta}<\frac{c_{1}c_{2}\rho_{1}}{96C_{1}}.$ (5.54)
Denoting $V=\text{Lie}(\text{SL}_{2})(\mathbb{Z})$. Multiplying LHS of (5.52) with its $g_{1}$ and $g_{2}$ conjugates, we obtain
Lemma 5.55.
There are $q_{5}||q$, $q_{5}>q^{\frac{c_{1}}{32C_{1}}}$, $q_{5}^{\prime}|q_{5}^{\prime\prime}|q_{5}$, and some constant $K_{2}>0$ depending only on $c_{1}$, such that
$1+q_{5}^{\prime}\mathbb{Z}V\subset A_{1}^{K_{2}}(\operatorname{mod}q_{5}^{\prime\prime}),$ (5.56)
where for each $p^{n}||q_{5}$, $v_{p}(q_{5}^{\prime\prime})=O_{c_{1}}(\rho_{2}n)$ and
$\frac{c_{1}c_{3}}{384C_{1}}\rho_{2}n<v_{p}(q_{5}^{\prime\prime}/q_{5})<v_{p}(q_{5}^{\prime\prime})/2$
if
$\rho_{1}<\frac{c_{1}c_{3}}{3\cdot 2^{9}C_{1}}\rho_{2}.$ (5.57)
5.6. Proof of Proposition 5.1
Collecting all the restrictions (5.5), (5.10), (5.19), (5.29), (5.30), (5.39), (5.50), (5.53), (5.54), (5.57) on $\rho_{1},\rho_{2},\delta$, we find we can take
$\rho_{1}=\frac{c_{1}^{2}\cdot c_{3}}{2^{8}3^{4}C_{1}}\rho_{2}$ (5.58)
$\delta=\frac{c_{1}c_{2}c_{3}}{96C_{1}}\rho_{1}$ (5.59)
We need the elementary lemma whose proof we leave as an exercise:
Lemma 5.60.
Let $H_{1},H_{2}\subset\text{SL}_{2}(\mathbb{Z})$ and let $p$ be a prime. Suppose $1+p^{m_{1}}V(\operatorname{mod}p^{m_{2}})\subset H_{1}(\operatorname{mod}p^{m_{2}})$ and $1+p^{n_{2}}V(\operatorname{mod}p^{n_{2}})\subset H_{2}(\operatorname{mod}p^{n_{2}})$ with $1\leq m_{1}\leq m_{2}\leq 2m_{1}$ and $1\leq n_{1}\leq n_{2}\leq 2n_{1}$. Then
$1+p^{m_{1}+n_{1}}V\subset(H_{1}H_{2})^{4}(\operatorname{mod}p^{m_{2}+n_{2}}).$
Proposition 5.1 is thus proved by iteratively applying Lemma 5.60 to (5.56), with $q_{*}=q_{5}$, $c=\frac{c_{1}}{32C_{1}}$ and $\rho=O_{c_{1}}(\rho_{2})$, $C=O_{S}(\log\frac{1}{\delta})$.
6. A gluing technique for the $\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ case
In this section and the next, we prove Proposition 2.2 for the case $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$. Suppose all the assumptions about $A$ hold but the conclusion (2.4) does not hold, then by taking $\delta$ sufficiently small, we will find a constant $K$ independent of $\delta$, such that $|\pi_{q}(A^{K})|>|\Lambda_{q}|^{1-\frac{\epsilon}{2}}$. But this will contradict Helfgott’s inequality (5.2) when $\delta$ is sufficiently small.
Proposition 5.1 so far only covers a large subset in one simple factor of the group and for some not too small modulus $q_{*}$. Apply Proposition 5.1 again with $q$ replaced by $q/q_{*}$, we will cover a large subset for some modulus $q_{*}^{\prime}$ coprime to $q_{*}$. This leads to considering the following question: Given a set $B\subset\Gamma_{q_{1}q_{2}},\text{gcd}(q_{1},q_{2})=1$, such that $\pi_{q_{1}}(B),\pi_{q_{2}}(B)$ are large in $\Gamma_{q_{1}},\Gamma_{q_{2}}$, is a bounded power of $B$ covers a large subset of $\Gamma_{q_{1}q_{2}}$?
The following proposition is the main gluing tool.

### v4

@@ 5.4. Constructing a line
$\zeta^{\prime}$ $\equiv 1+\zeta h_{1}-h_{1}\zeta\hskip 5.69054pt(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}\mathcal{\langle}\mathscr{Q}_{1},\mathscr{Q}_{2}\rangle)$
$\equiv 1+(\zeta-1)h_{1}-h_{1}(\zeta-1)\hskip 5.69054pt(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}\mathcal{\langle}\mathscr{Q}_{1},\mathscr{Q}_{2}\rangle)$
$\Rightarrow\hskip 8.5359pt$ $\zeta^{\prime}\equiv 1+\mathfrak{q}_{1}\mathfrak{q}_{2}\left(\begin{matrix}0&-\lambda_{h_{1}}Y\\ \lambda_{h_{1}}Z&0\end{matrix}\right)(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}\mathscr{Q}_{3}),$ (5.69)
Next, applying Lemma 5.67 with $x=\zeta^{\prime}$, $\mathcal{I}_{1}=\mathscr{Q}_{1}\mathscr{Q}_{2}$, $y=h_{2}\in H$, $\mathcal{I}_{2}=\mathscr{Q}_{2}$, we obtain
$\zeta^{\prime}h_{2}\zeta^{\prime-1}h_{2}^{-1}\equiv 1(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}^{2})$
and
$\zeta^{\prime}h_{2}\zeta^{\prime-1}h_{2}^{-1}\equiv 1+\zeta^{\prime}h_{2}-h_{2}\zeta^{\prime}\equiv 1+(\zeta^{\prime}-1)h_{2}-h_{2}(\zeta^{\prime}-1)(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}^{2}\mathscr{Q}_{3}),$ (5.70)
where we notice that $\mathscr{Q}_{3}|\mathscr{Q}_{2}$.
As $\zeta^{\prime}-1\equiv 0(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2})$ and $h_{2}\equiv\left(\begin{matrix}a_{h_{2}}&0\\ 0&a_{h_{2}}^{-1}\end{matrix}\right)(\operatorname{mod}\mathscr{Q}_{2}\mathscr{Q}_{3})$, we have
$(\zeta^{\prime}-1)h_{2}-h_{2}(\zeta^{\prime}-1)$ $\equiv(\zeta^{\prime}-1)\left(\begin{matrix}a_{h_{2}}&0\\ 0&a_{h_{2}}^{-1}\end{matrix}\right)-\left(\begin{matrix}a_{h_{2}}&0\\ 0&a_{h_{2}}^{-1}\end{matrix}\right)(\zeta^{\prime}-1)(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}^{2}\mathscr{Q}_{3})$ (5.71)
$\equiv\mathfrak{q}_{1}\mathfrak{q}_{2}^{2}\left(\begin{matrix}0&\lambda_{h_{1}}\lambda_{h_{2}}Y\\ \lambda_{h_{1}}\lambda_{h_{2}}Z&0\end{matrix}\right)(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}^{2}\mathscr{Q}_{3}),$ (5.72)
(5.70) and (5.71) then gives
$\zeta^{\prime}h_{2}\zeta^{\prime-1}h_{2}^{-1}\equiv 1+\mathfrak{q}_{1}\mathfrak{q}_{2}^{2}\left(\begin{matrix}0&\lambda_{h_{1}}\lambda_{h_{2}}Y\\ \lambda_{h_{1}}\lambda_{h_{2}}Z&0\end{matrix}\right)(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}^{2}\mathscr{Q}_{3}).$
Applying Lemma 5.67 iteratively, we have for any $h_{1},h_{2},\cdots,h_{C_{3}}\in\mathscr{H}$,
$[\cdots[[\zeta,h_{1}],h_{2}],\cdots h_{C_{3}}]\equiv 1(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}^{C_{3}}),$ (5.73)
$[\cdots[[\zeta,h_{1}],h_{2}],\cdots h_{C_{3}}]\equiv 1+\mathfrak{q}_{1}\mathfrak{q}_{2}^{C_{3}}\left(\begin{matrix}0&(\prod_{i=1}^{C_{3}}\lambda_{h_{i}})(-1)^{C_{3}}\cdot Y\\ (\prod_{i=1}^{C_{3}}\lambda_{h_{i}})Z&0\end{matrix}\right)(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}^{C_{3}}\mathscr{Q}_{3})$ (5.74)
The sum structure is clear. For any $h_{1,1},\cdots,h_{C_{2},C_{3}}\in\mathscr{H}$, write
$u_{j}=[\cdots[[\zeta,h_{j,1}],h_{j,2}],\cdots h_{j,C_{3}}],\hskip 14.22636pt1\leq j\leq C_{2}.$
We have
$u_{1}u_{2}\cdots u_{C_{2}}\equiv 1+\mathfrak{q}_{1}\mathfrak{q}_{2}^{C_{3}}\sum_{i=1}^{C_{2}}\left(\prod_{j=1}^{C_{3}}h_{\sigma_{i,j}}\right)\left(\begin{matrix}0&(-1)^{C_{3}}Y\\ Z&0\end{matrix}\right)(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}^{C_{3}}\mathscr{Q}_{3})$ (5.75)
From (5.46) and (5.75), we conclude there is a constant $K_{1}$ depending only on $S$ such that
$1+\mathfrak{q}_{1}\mathfrak{q}_{2}^{C_{3}}\mathbb{Z}\xi\left(\begin{matrix}0&(-1)^{C_{3}}Y\\ Z&0\end{matrix}\right)\subset M^{-1}A_{0}^{K_{1}}M(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}^{C_{3}}\mathcal{Q}^{\prime})$ (5.76)
where $\mathcal{Q}^{\prime}=\langle\mathcal{Q},\mathscr{Q}_{3}\rangle$ with $\mathcal{Q}$ given in (5.46).
Now for each $p^{n}\|q_{4}$, we measure the thickness of the above progression localized at $\mathcal{P}_{p}$ in terms of the valuation $v_{\mathcal{P}_{p}}$. We have
[…]
where for the second equality we have used that $\mathcal{Q}$ divides $(q_{4}^{\{w_{2}\}})$ and $\mathscr{Q}_{3}=(q_{4}^{\{w_{2}-O(\rho_{1})\}})$. From the definition of $q_{2}$ (5.48), we have
$v_{\mathcal{P}_{p}}(\mathcal{Q})-v_{\mathcal{P}_{p}}(\xi)>\frac{nw_{2}\varepsilon_{0}}{4},$
and so
$v_{\mathcal{P}_{p}}(\mathscr{Q}_{1}\mathscr{Q}_{2}^{C_{3}}\mathcal{Q}^{\prime})-v_{\mathcal{P}_{p}}(\mathfrak{q}_{1}\mathfrak{q}_{2}^{C_{3}}\xi Z)>\frac{nw_{2}\varepsilon_{0}}{5}$ (5.77)
by taking $L_{2}$ large.
Collecting (5.76) and (5.77), and conjugating (5.76) back by $M$, we obtain
@@
There is a constant $K_{1}$ depending only on $S$ and $\varepsilon$, an exact divisor $q_{4}\|q_{2}$, $q_{4}\geq q^{\Omega(1)}$, an ideal $\mathcal{U}\subset\mathcal{O}$ dividing $(q_{4})$, an element $\mathfrak{u}\in\mathcal{O}$, and a primitive traceless matrix $W\in\text{ Mat}_{2}(\mathcal{O})$, such that
$1+\mathbb{Z}\mathfrak{u}W\subset A_{0}^{K_{1}}(\operatorname{mod}\mathcal{U}),$ (5.79)
where for each $p^{n}\|q_{4}$, we have
$v_{\mathcal{P}_{p}}(\mathcal{U})=\Theta(\rho_{2}n),$ (5.80)
$\frac{v_{\mathcal{P}_{p}}(\mathcal{U})}{2}\geq v_{\mathcal{P}_{p}}(\mathcal{U})-v_{\mathcal{P}_{p}}(\mathfrak{u})\geq\Omega(\rho_{2}n).$ (5.81)
Proof of Lemma 5.50.
Let $r_{m}\in A_{0}^{K_{1}},m\in\mathbb{Z}$ such that
$r_{m}\equiv 1+m\mathfrak{u}W(\operatorname{mod}\mathcal{U})$ (5.82)
as given in Lemma 5.78. Write
$r_{1}\equiv 1+Q_{1}X(\operatorname{mod}Q_{1}^{2}).$
where $Q_{1}|q_{4}$ and $X\in\text{Mat}_{2}(\mathbb{Z})$ is traceless and coprime to $q_{4}$. The choice for $Q_{1}$ is unique.
For each $p^{n}\|q_{4}$, we have
$v_{\mathcal{P}_{p}}(Q_{1})=v_{\mathcal{P}_{p}}(\mathfrak{u}).$ (5.83)
and so
$v_{p}(Q_{1})=\frac{v_{\mathcal{P}_{p}}(\mathfrak{u})}{v_{\mathcal{P}_{p}}(p)}.$ (5.84)
It follows from (5.82) that
$\forall m\in\mathbb{Z},\hskip 14.22636ptr_{1}^{m}r_{m}^{-1}\equiv 1(\operatorname{mod}\mathcal{U}).$ (5.85)
Taking Galois conjugate of (5.85), we obtain
$\forall m\in\mathbb{Z},\hskip 14.22636ptr_{1}^{m}r_{m}^{-1}\equiv 1(\operatorname{mod}\bar{\mathcal{U}})$ (5.86)
Let $\mathcal{U}\cap\bar{\mathcal{U}}=(Q_{2})$ for some $Q_{2}\in\mathbb{N}$, $Q_{2}|q_{4}$. Then (5.85) and (5.86) gives
$\forall m\in\mathbb{Z},\hskip 14.22636ptr_{1}^{m}r_{m}^{-1}\equiv 1(\operatorname{mod}Q_{2})$ (5.87)
by Chinese Remainder Theorem, which implies
$r_{m}\equiv 1+mQ_{1}X(\operatorname{mod}Q_{2}).$ (5.88)
We also have $v_{\mathcal{P}_{p}}(Q_{2})=v_{\mathcal{P}_{p}}(\mathcal{U})$, which can be seen from the prime factorization of $\mathcal{U}$, so
$v_{p}(Q_{2})=\frac{v_{\mathcal{P}_{p}}(Q_{2})}{v_{\mathcal{P}_{p}}(p)}=\frac{v_{\mathcal{P}_{p}}(\mathcal{U})}{v_{\mathcal{P}_{p}}(p)}.$ (5.89)
(5.84) and (5.89) give
$v_{p}(Q_{2})-v_{p}(Q_{1})=\frac{v_{\mathcal{P}_{p}}(\mathcal{U})-v_{\mathcal{P}_{p}}(\mathfrak{u})}{v_{\mathcal{P}_{p}}(p)}.$ (5.90)
(5.90), (5.81), (5.84), (5.89) and (5.90) then give the desired estimates for valuations in Lemma 5.50.
5.5. Constructing a segment
We want to find $g_{1},g_{2}$ from a bounded product of $\mathbb{P}_{1}(A)$ so that $g_{1},g_{2}$ conjugates the arithmetic progressions given in Lemma (5.50) to different directions. Multiplying them together then creates a thick segment. Lemma 5.92 is our goal.
Identify $\text{Lie(SL}_{2})(\mathbb{Z})$ with $\mathbb{Z}^{3}$ by $\left(\begin{matrix}a&b\\ c&-a\end{matrix}\right)\mapsto(a,b,c)^{t}$. Under this identification, write $X_{0}=(x_{1},x_{2},x_{3})$.
Choose an element $T_{1}\in\text{SL}_{3}(\mathbb{Z})$ so that $T_{1}(x_{1},x_{2},x_{3})^{t}=(1,0,0)^{t}$. Define $\mathcal{L}_{1}:\mathbb{Z}^{3}\rightarrow\mathbb{Z}$ by $\mathcal{L}_{1}(a,b,c)=b$. Then $\mathcal{L}_{1}\circ T_{1}$ is a primitive linear form on $\mathbb{Z}^{3}$.
We apply Proposition 4.2 with $\xi_{1}=X_{0},\xi_{2}=\eta_{2}=0$, and a proper choice of $\eta_{1}$ so that
$\text{Tr}(gX_{0}g^{-1}\eta_{1})=\mathcal{L}_{1}\circ T_{1}(gX_{0}g^{-1}).$
Recall $c_{2}$ be the implied constant from Propositions 4.2. Then we obtain, for any sufficiently large $Q>q_{4}^{\rho_{1}}$,
$\nu^{(l)}(\{g\in\text{SL}_{2}(\mathbb{Z})|\mathcal{L}_{1}\circ T_{1}(gX_{0}g^{-1})\equiv 0(\operatorname{mod}{Q})\})<Q^{-c_{2}},$
where $l$ is as given in Proposition 2.2. This implies, arguing in the same way as for the construction of $A_{1}$ in §5, that there is $q_{5}\|q_{4}$, $q_{5}>(q_{4})^{\frac{1}{2}}$, and $g_{1}\in\mathbb{P}_{1}(A)$ such that for any $p^{n}\|q_{5}$,
$\mathcal{L}_{1}\circ T_{1}(g_{1}Xg_{1}^{-1})\neq 0(p^{[2\rho_{1}n]}),$
by taking $L_{1}$ at (5.5) sufficiently large.
Let $X_{1}=g_{1}X_{0}g_{1}^{-1}$, then we can write
$X_{1}\equiv a(1,0,0)+b\vec{v}(\operatorname{mod}q_{5}),$
where $a,b\in\mathbb{Z}$ with $b$ satisfying, for any $p^{n}\|q_{5}$, $p^{[2\rho_{1}n]}\nmid b$, and $\vec{v}$ is a primitive vector in the $\mathbb{Z}$-span of $(0,1,0)$ and $(0,0,1)$.
Next, we choose an element $T_{2}\in\text{SL}_{3}(\mathbb{Z})$ such that $T_{2}(X_{0})=(1,0,0)$ and $T_{2}(X_{1})=(0,b,0)$. Let $\mathcal{L}_{2}:\mathbb{Z}^{3}\rightarrow\mathbb{Z}$ be the linear form getting the third component. Then $\mathcal{L}_{2}\circ T_{2}$ is a primitive linear form on $\mathbb{Z}^{3}$. Applying Proposition 4.2 again, for any $Q>q_{5}^{\rho_{1}}$, we have
$\nu^{(l)}(\{g\in\text{SL}_{2}(\mathbb{Z})|\mathcal{L}_{2}\circ T_{2}(gX_{0}g^{-1})\equiv 0(\operatorname{mod}{Q})\})<Q^{-c_{2}},$
which then implies there is $q_{6}\|q_{5}$, $q_{6}>(q_{5})^{\frac{1}{2}}$, and $g_{2}\in\mathbb{P}_{1}(A)$ such that
$c:=\mathcal{L}_{2}\circ T_{2}(g_{2}Xg_{2}^{-1})\not\equiv 0(p^{[2\rho_{1}n]})$
for any $p^{n}\|q_{6}$. Therefore, for any $p^{n}\|q_{6}$,
$\text{Det}(X_{0},X_{1},X_{2})=\text{Det}\left(\begin{matrix}1&0&0\\ 0&b&0\\ *&*&c\end{matrix}\right)\not\equiv 0(\operatorname{mod}p^{[4\rho_{1}n]}).$ (5.91)
Denoting $V=\text{Lie}(\text{SL}_{2})(\mathbb{Z})$. Multiplying LHS of (5.50) with its $g_{1}$ and $g_{2}$ conjugates, with (5.91) in mind, we obtain
Lemma 5.92.
There is an exact divisor $q_{6}\|q_{4}$, $q_{6}\geq q_{4}^{1/4}$, $Q_{3}|Q_{4}|q_{6}$, and some constant $K_{2}>0$ depending only on $S$ and $\varepsilon$, such that
$1+Q_{3}V(\operatorname{mod}Q_{4})\subset A_{0}^{K_{2}}(\operatorname{mod}Q_{4}),$ (5.93)
where for each $p^{n}\|q_{6}$,
$v_{p}(Q_{4})=\Theta(\rho_{2}n),$ (5.94)
$v_{p}(Q_{4})-v_{p}(Q_{3})=\Theta(\rho_{2}n).$ (5.95)
5.6. Proof of Proposition 5.1
We iteratively apply Lemma 3.25 to sets starting with $E\subset A_{0}^{K_{2}}$ implied by (5.93), until we create a set $E_{0}$ such that
$1+Q_{5}V(\operatorname{mod}Q_{6})\subset E_{0}(\operatorname{mod}Q_{6}),$
where for each $p^{n}\|q_{6}$, $v_{p}({Q_{6}})=\Theta(\rho_{2}n)$ and
$v_{p}{(Q_{6})}-v_{p}{(Q_{5})}\geq v_{p}(Q_{3}).$ (5.96)
Then we iteratively taking commutators and obtain
$E_{1}=[E_{0},E],1+Q_{3}Q_{5}V(\operatorname{mod}Q_{3}Q_{6})\subset E_{1}(\operatorname{mod}Q_{3}Q_{6});$
$\cdots$
$E_{k}=[E_{k-1},E],1+(Q_{3})^{k}Q_{5}V(\operatorname{mod}(Q_{3})^{k}Q_{6})\subset E_{k}(\operatorname{mod}(Q_{3})^{k}Q_{6});$
$\cdots$
until we reach a set with all elements congruent to identity mod $q_{6}$, which takes $O(\frac{1}{\rho_{2}})$ steps. Since
@@ 5.5. Constructing a segment
we deduce
$E_{0}E_{1}E_{2}\cdots(\operatorname{mod}q_{6})\supset\Gamma(Q_{5})/\Gamma(q_{6}).$
Proposition 5.1 is thus proved with $q_{*}=q_{6}$, $c=O(1)$, $\rho=\Theta(\rho_{2})$ and $C=O(\frac{1}{\rho_{2}})$.
6. A gluing technique for $\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$
In this section, we continue our progress towards proving Proposition 2.2. The set $S$ and the constant $\varepsilon>0$ in Proposition 2.2 are fixed. We still let $c_{1}$ and $c_{2}$ be the implied constants for $c$ in Propositions 4.1 and 4.2 respectively. We let $c_{0}$ be the implied constant for $c$ in Proposition 5.1. We free all other notations, such as $\mathbb{P}_{1},\mathbb{P}_{2},q_{1},q_{2}\cdots$ and $Q_{1},Q_{2}\cdots$ newly introduced in Section 5. The Landau notations $O,\Omega,\Theta$ describe asymptotic behaviours of functions with implied constants depending only on $S$.
Proposition 5.1 so far only covers a very large subset for some not too small modulus $q_{*}$ in one simple factor of the group. For the purpose of proving Proposition 2.2, we need to grow $q_{*}$ to a very large divisor of $q$, and also cross from one simple factor to another simple factor. Our main goal in this section is to prove the following gluing tool: